# GaussianEditor: Editing 3D Gaussians Delicately with Text Instructions

### v1

@@ 3.5 Delicate Editing within Gaussian RoI
To achieve delicate editing in 3D scenes, we use the Gaussian RoI to constrain the editing area. In particular, we randomly sample viewpoints from the 3D scene representation and render 2D image representations ${\mathcal{I}}_{rd}$. Then 2D diffusion model ${\mathcal{D}}$ is used to edit the rendered image ${\mathcal{I}}_{rd}$, with the user instruction ${\mathcal{T}}$ and original image ${\mathcal{I}}_{ori}$ as conditions, to get the edited image ${\mathcal{I}}_{ed}$ as
${\mathcal{I}}_{ed}={\mathcal{D}}({\mathcal{I}}_{rd},t;{\mathcal{T}},{\mathcal{I}}_{ori}),$ (12)
@@
where ${\mathcal{R}}$ is the set of RoI attributes. That means only Gaussians in RoI can receive gradients. Finally, we utilize the Adam algorithm to optimize the 3D Gaussians. After $it$ rounds of training, the edited 3D scene ${\mathcal{G}}_{ed}$ is obtained.
@@ 4.2 Qualitative Evaluation
In Fig. 1 and Fig. 4, we present the visual results of our method, demonstrating the precise editing effects while ensuring 3D consistency. Fig. 1 demonstrates the editing capabilities of our method on human subjects. The first column displays the original scenes. In the second column, the first row “Give him a red nose” illustrates our ability to change colors. The third row in the second column, “Make him completely bald”, showcases our geometric editing capability. The second row in the second column demonstrates our precise editing ability by exclusively editing the left side of the person’s face. Building upon this foundation, we achieve editing in the third column, specifically focusing on the right side of the face, showcasing our ability to perform multi-round edits and accurately fulfill user instructions. Fig. 4 showcases the precise editing capabilities of our method in open 3D scenes. In the upper portion, the bicycle scene allows us to accurately locate the position of the road and edit its texture, transforming it into the grass, a river. It is worth noting that in the experiment where we change the texture to a river, our method accurately constructs the reflection, making it appear realistic. On the basis of editing the road into a river, we further edited the bench, proving that our method can achieve multiple rounds of editing. The lower portion demonstrates the results of editing the bear using our method, where we fully preserve the original appearance of the background area and focus the edits on the bear.
@@
Fig. 5 compares the results of our method with those of IN2N [11], on the scenes presented in IN2N. From the figure, it is evident that our method successfully changes the texture of pants without affecting the clothes, and vice versa, demonstrating the effectiveness of our method in distinguishing different objects within the foreground. Additionally, when editing the clothes and pants, the background remains unaffected, indicating our method’s effective separation of foreground and background. Furthermore, the last column reveals that IN2N, limited by 2D diffusion, distorts the face, while our method maintains a superior rendering quality of faces.
Figure 5: Comparisons with Instruct-NeRF2NeRF (IN2N) [11]. We use IN2N as a baseline and compare our method with the scenes presented in their paper.
@@
Furthermore, we present the results of our editing in a complex scene featuring multiple objects, as depicted in Fig 6. We specifically selected three distinct object types for editing purposes. The first object type is the background. In a particular sub-scene, the desktop shares similarities with the background. Employing a caption-based approach, we successfully transformed the desktop into a wooden material. As evident in the image, the edited result exhibits a distinct and pronounced wood texture. The second object type is a foreground object. In the sub-scene, the flowerpot serves as one of the foreground objects. We opted to change the color of the flowerpot to red, and the outcome was highly successful. Lastly, the most intricate editing task involved the rolling pin, which was occluded by multiple objects from various perspectives. As illustrated in the lower right corner of the picture, the rolling pin was hindered by different objects, yet we managed to edit it into a cucumber without any adverse impact on the other objects.
@@ 4.3 Ablation Study and Analysis
To validate the effectiveness of each module in our framework, we design three variant approaches: (1) w/o Gaussian RoI: We discontinued the use of Gaussian RoI ${\mathcal{G}}_{RoI}$ to control the gradients of Gaussian points, as mentioned in Eq. 14. (2) w/o Text ROI: In this scenario, we ceased the selection of text ROI ${\mathcal{T}}_{RoI}$ using LLM assistant ${\mathcal{A}}$. Instead, all the words in the user’s instruction are treated as text RoI ${\mathcal{T}}_{RoI}$. (3) w/o RoI lifting: Instead of lifting the image RoI ${\mathcal{I}}_{RoI}$ to 3D Gaussians, the image RoI ${\mathcal{I}}_{RoI}$ is used to govern the calculation of the loss. That is, only the pixels within the image RoI ${\mathcal{I}}_{RoI}$ are taken into account for the loss computation. Fig. 7 showcases the outcomes of our ablation experiment, which aimed to edit the doll based on the instruction “Turn its mouth red.” The results reveal the following findings. When the Gaussian RoI is not used, the 3D scene is all turned red because the 2D diffusion fails to successfully control the editing area. In cases where text ROI ${\mathcal{T}}_{RoI}$ is not utilized, the grounding segmentation model tends to segment the entire foreground object, leading to the doll being entirely edited to red. (3) When RoI lifting ${\mathcal{H}}$ is not employed, the doll’s mouth is successfully turned red, but other facial areas are also affected. We also find that the current 2D grounding segmentation model may fail to parse specific views, which means there exist noises on the image RoI ${\mathcal{I}}_{RoI}$. Consequently, when optimizing the scene based on the ${\mathcal{I}}_{RoI}$, leakage occurs during the editing process. Our proposed RoI lifting module effectively addresses this issue during training. In conclusion, our ablation experiment effectively demonstrates the effectiveness of several RoI-related modules in our method.
Figure 7: Ablation results of settings without Gaussian RoI, without Text RoI, and without RoI Lifting.
@@
We further conduct experiments to evaluate the role of scene description generation, employing three distinct experimental setups. The first one solely relies on the utilization of the template ${\mathcal{T}}_{tmp}$ and user instruction ${\mathcal{T}}$ to compose the user message ${\mathcal{T}}_{usr}$, without employing scene description generation. The LLM was employed to acquire the text RoI ${\mathcal{T}}_{RoI}$. The second one involves randomly sampling a view and extracting the corresponding image’s text description ${\mathcal{T}}_{img}$, which was then incorporated into the template ${\mathcal{T}}_{tmp}$ as the text description ${\mathcal{T}}_{dsc}$ and user instructions ${\mathcal{T}}$ to form the user message ${\mathcal{T}}_{usr}$, which can be denoted as ${\mathcal{T}}_{usr}={\mathcal{T}}_{tmp}({\mathcal{T}}_{img},{\mathcal{T}})$. The third one represents the complete version of our approach. The test scene involves a park where a bike and a bench are positioned closely together. The editing instruction is “Turn the thing next to the bike orange”. The obtained results are presented in Fig. 8. As is evident from the image, when description generation is not employed, the LLM fails to acquire the Text ROI according to the user instructions, resulting in editing failure. The second one randomly samples images to obtain text descriptions, resulting in incomplete descriptions and leading to an incorrect text ROI prediction by the LLM. Consequently, the final editing result turns the road into an orange color. In contrast, our method flawlessly executes the editing task. This success can be attributed to scene description generation, which obtains an accurate text description encompassing the relative positional relationship between the bicycle and the bench. This enables the LLM to analyze and determine the user’s intention to edit the bench. Consequently, the desired color change of the bench is successfully implemented.
@@ 4.4 Limitations
Although our framework has solved some problems inherited from the integrated sub-modules, e.g. noise in the results of grounding segmentation, there are still some problems that the current system cannot completely avoid. In scene description generation, the descriptions from different views of the same object may differ from each other. When the differences are large enough, the LLM may misunderstand these descriptions as those from multiple objects. This issue does not affect the results in the current experiment, but we would like to optimize this in the future. In addition, our system cannot achieve good editing results in scenes where the grounding segmentation or diffusion model completely fails.
@@ 5 Conclusion
This paper proposes a systematic framework, named GaussianEditor, for text-guided delicate 3D scene editing. As we know, GaussianEditor is the first work to edit 3D Gaussians, taking advantage of the explicit property of 3D Gaussians and making it easy to control the editing area precisely. Several techniques are proposed to achieve delicate editing, including extracting instruction RoI from texts, aligning the RoI to 3D Gaussians, and editing the scene with the Gaussian RoI. GaussianEditor achieves notably more delicate editing results than IN2N [11] with much shorter training time (within 20 minutes v.s. 45 minutes – 2 hours). Noticing recent works [27, 50] have extended Gaussian splatting to dynamic scenes, we leave the delicate editing in dynamic scenes as future work.
@@ A.2 Quantitative Evaluation
In Tab. 1, we present the quantitative evaluation results. We follow the metrics used in Instruct NeRF2NeRF (IN2N) [11], including the CLIP [36] text-image direction similarity and image-image similarity between the original scene and the edited scene. The quantitative results indicate that our method achieves a comparable CLIP text-image direction similarity score with IN2N, while image-image similarity has improved a lot. We would like to analyze the limitations of the used metric as follows.
Table 1: Results of CLIP Text-Image Direction Similarity and Image-Image Similarity between the original scene and edited scene.
@@
A.3 Extension

### v2

@@ 3.5 Delicate Editing within Gaussian RoI
To achieve delicate editing in 3D scenes, we use the Gaussian RoI to constrain the editing area. In particular, we randomly sample viewpoints from the 3D scene and render 2D image ${\mathcal{I}}_{render}$. After that, ${\mathcal{I}}_{render}$ and noise level $t$ are put into 2D diffusion model ${\mathcal{M}}_{DM}$, with the user instruction ${\mathcal{T}}$ and image ${\mathcal{I}}_{input}$ of input scene as conditions, to get edited image ${\mathcal{I}}_{edit}$:
${\mathcal{I}}_{edit}={\mathcal{D}}({\mathcal{I}}_{render},t;{\mathcal{T}},{\mathcal{I}}_{input}),$ (12)
@@
where ${\mathcal{R}}$ is the set of RoI attributes. That means only Gaussians in RoI can receive gradients. Finally, we utilize the Adam algorithm to optimize the 3D Gaussians. After many rounds of training, the edited 3D scene ${\mathcal{G}}_{edit}$ is obtained.
@@ 4.2 Qualitative Evaluation
In Fig. 1 and Fig. 4, we present the visual results of GaussianEditor, demonstrating the precise editing effects while ensuring 3D consistency. Fig. 1 shows the editing capabilities for characters. The first column displays the original scenes. In the second column, the first row “Give him a red nose” illustrates color-changing ability, while the third row, “Make him completely bald”, showcases capabilities of retexturing and slight geometry editing. The second row in the second column demonstrates precise editing ability by exclusively editing the left side of the face. Based on that, we achieve editing in the third column, focusing on the right side of the face, showcasing the ability of multi-round edits, and accurately fulfilling user instructions. Fig. 4 showcases the precise editing capabilities in open 3D scenes. In the upper portion, the bicycle scene allows us to accurately locate the position of the road and edit its texture, transforming it into the grass, a river. In the experiment where we change the texture to a river, our method accurately constructs the reflection, making it appear realistic. Based on editing the road into a river, we further edited the bench, proving that our method can achieve multiple rounds of editing. The lower portion demonstrates the results of editing the bear, which fully preserves the original appearance of the background area and focuses the edits on the bear.
@@
Fig. 5 compares the results of our method with those of IN2N [11], on the scenes presented in IN2N. From the figure, it is evident that our method changes the texture of the pants without affecting the clothes, and vice versa, demonstrating the effectiveness of our method in distinguishing different objects within the foreground. Additionally, when editing the clothes and pants, the background remains unaffected, indicating our method’s effective separation of foreground and background. Furthermore, the last column reveals that IN2N, limited by 2D diffusion, distorts the face, while our method maintains a superior rendering quality of faces.
Figure 5: Comparisons with Instruct-NeRF2NeRF (IN2N) [11] on the scene presented in their paper.
@@
Furthermore, we present the results of our editing in a complex scene featuring multiple objects, as depicted in Fig 6. Three distinct object types are selected for editing purposes. The first type is the background, which is the desktop in this scene. We successfully transformed the desktop into a wooden material using a caption-based approach. The edited result exhibits a distinct wood texture. The second object type is a foreground object, the flowerpot. We opted to change the color of the flowerpot to red, and the outcome was highly successful. Lastly, the most intricate editing task involved the rolling pin, which was occluded by multiple objects from various perspectives. As shown in the lower right corner of the picture, we managed to edit it into a cucumber without impacting the other objects.
Table 1: Quantitative evaluation on the bicycle scene of the Mip-NeRF360 dataset [3].
[…]
IN2N [11] 0.22 0.85 103 51 min
Ours-DVGO [46] 0.11 0.82 148 40 min
Ours-3DGS 0.28 0.95 51 20 min
4.3 Quantitative Evaluation
Metric Comparisons.
Table 1 shows quantitative results on the bicycle scene of the Mip-NeRF360 dataset [3], comparing with IN2N [11] and Direct Voxel Grid Optimization (DVGO) [46] as the representation. The metrics include CLIP Text-Image Direction Similarity (CTIDS), Image-Image Similarity (IIS), FID, and training time. GaussianEditor achieves the best results in all metrics. The test data is shown in the supplementary material.
User Study.
We perform a user study comparing with IN2N on the bear scene in Fig. 4 and the human scene in Fig. 5, involving 21 participants. GaussianEditor gets a 87.07% voting percentage, while IN2N gets 12.93%.
@@ 4.4 Ablation Study and Analysis
To validate the effectiveness of each module in our framework, we design three variant approaches: (1) w/o Gaussian RoI: We discontinued the use of Gaussian RoI ${\mathcal{G}}_{RoI}$ to control the gradients of Gaussian points, as mentioned in Eq. 14. (2) w/o Text ROI: In this scenario, we ceased the selection of text ROI ${\mathcal{T}}_{RoI}$ using LLM assistant ${\mathcal{M}}_{LLM}$. Instead, all the words in the user’s instruction are put to ${\mathcal{M}}_{Seg}$ to get ${\mathcal{I}}_{RoI}$. (3) w/o RoI lifting: Instead of lifting the image RoI ${\mathcal{I}}_{RoI}$ to 3D Gaussians, the image RoI ${\mathcal{I}}_{RoI}$ is used to govern the calculation of the loss. That is, only the pixels within the image RoI ${\mathcal{I}}_{RoI}$ are taken into account for the loss computation. Fig. 7 showcases the outcomes of our ablation experiment, which aimed to edit the doll based on the instruction “Turn its mouth into red.” The results reveal the following findings. (1) When the Gaussian RoI is not used, the 3D scene is all turned red because the 2D diffusion fails to control the editing area. (2) In cases where text ROI ${\mathcal{T}}_{RoI}$ is not utilized, the grounding segmentation model tends to segment the entire foreground object, leading to the doll being entirely edited to red. (3) When RoI lifting ${\mathcal{M}}_{Lift}$ is not employed, the doll’s mouth is successfully turned red, but other facial areas are also affected. Because the grounding segmentation model may fail to parse specific views, noise exists on the image RoI ${\mathcal{I}}_{RoI}$. Consequently, leakage occurs during the editing process. Our proposed RoI lifting module effectively addresses this issue during training. In conclusion, our ablation experiment demonstrates the effectiveness of several RoI-related modules in our method.
Figure 7: Ablation experiment of RoI.
@@
We further conduct experiments to evaluate the role of scene description generation, employing three distinct experimental setups. The first one composes the user message ${\mathcal{T}}_{user}$, without employing scene description. The second method randomly samples a view and extracts the corresponding image’s text description as the scene description. The third one represents the complete version of our approach. The test scene involves a park where a bike and a bench are positioned closely together. The editing instruction is “Turn the thing next to the bike orange”. The obtained results are presented in Fig. 8. As shown in the image, when scene description is not employed, the LLM fails to acquire the Text ROI according to the user instructions, resulting in editing failure. The second one randomly samples images to obtain scene descriptions, resulting in incomplete descriptions and leading to an incorrect text ROI prediction by the LLM. Consequently, the final editing result turns the road into an orange color. In contrast, our method flawlessly executes the editing task. This success can be attributed to scene description generation, which obtains an accurate text description encompassing the relative positional relationship between the bicycle and the bench. This enables the LLM to analyze and determine the user’s intention to edit the bench. Consequently, the desired color change of the bench is successfully implemented.
@@ 4.5 Limitations
Although our framework has solved some problems inherited from the integrated sub-modules, e.g. noise in the results of grounding segmentation, there are still some problems that the current system cannot completely avoid. In scene description generation, the descriptions from different views of the same object may differ from each other. When the differences are large enough, the LLM may misunderstand these descriptions as those from multiple objects. This issue does not affect the results in the current experiment, but we would like to optimize this in the future. In addition, our system cannot achieve good editing results in scenes where the grounding segmentation or diffusion model completely fails, such as drastic geometric editing.
@@ 5 Conclusion
This paper proposes a systematic framework, named GaussianEditor, for text-guided delicate 3D scene editing. As we know, GaussianEditor is one of the first works to edit 3D Gaussians, taking advantage of the explicit property of 3D Gaussians and making it easy to control the editing area precisely. Several techniques are proposed to achieve delicate editing, including extracting instruction RoI from texts, aligning the RoI to 3D Gaussians, and editing the scene with the Gaussian RoI. GaussianEditor achieves notably more delicate editing results than IN2N [11] with much shorter training time (within 20 minutes v.s. 45 minutes – 2 hours). Noticing recent works [27, 50, 55, 56] have extended Gaussian splatting to dynamic scenes, we leave the delicate editing in dynamic scenes as future work.
@@ A.2 Quantitative Evaluation
In Tab. 2, we present the quantitative evaluation results. The scenes in Fig. 5 are used for this test. We follow the metrics used in Instruct NeRF2NeRF (IN2N) [11], including the CLIP [36] text-image direction similarity and image-image similarity between the original scene and the edited scene. The quantitative results indicate that our method achieves a comparable CLIP text-image direction similarity score with IN2N, while image-image similarity has improved a lot. We would like to analyze the limitations of the used metric as follows.
Table 2: Results of CLIP Text-Image Direction Similarity and Image-Image Similarity between the original scene and edited scene. Test scene is shown in Fig. 5.
@@
Figure 11: Visualization result of Tab. 1.
User Study.
Here are more details of the user study shown in Sec. 4.3. 4 human editing results in Fig. 5 and 3 bear editing results in Fig. 4 are chosen for the user study, forming 7 questions for the questionnaire. In every question, we showcase the original scene, the text instructions for editing, and the editing results of IN2N [11] and GaussianEditor. For equality, the editing results in the question are randomly named using the letter A or B. Users are required to choose the better one. After 21 users submit their questionnaires, 147 votes (21 users $\times$ 7 questions) are collected. GaussianEditor gets 128 votes for all questions and IN2N gets 19 votes, accounting for 87.07% and 12.93%, respectively.
A.3 Qualitative Evaluation
Comparison with IN2N [11] and Different Backbones.
In Fig. 11, we show the qualitative result of IN2N and GaussianEditor with different backbones. This scene is also used in Tab. 1. IN2N fails in this task and turns the bicycle, bench, and tree all red. Besides, the backbone using DVGO [46] also has difficulty in localizing the bench precisely and produces worse rendering results, while GaussianEditor grounds the bench precisely and turns it red.
Comparison with DreamEditor [59].
In Fig. 12, we show the qualitative result of DreamEditor and GaussianEditor. GaussianEditor delicately edits the doll and retains the hair details, while DreamEditor wipes the hair and changes the back box. Besides, GaussianEditor gets the wanted editing result using less time.
Depth Map of Geometric Editing.
In Fig. 13, we show the depth map of the hair editing result shown in Fig.1. The depth map indicates that GaussianEditor possesses a certain level of geometric editing capability. The task of handling drastic geometric editing changes is left for future work.
Figure 12: Comparison to DreamEditor on DTU dataset.
Figure 13: Depth map of hair editing in Fig.1.
A.4 Extension